Linda is considering investing in a company's stock and is aware that the return on that investment is particularly sensitive to how the economy is performing. Her analysis suggests that four states of the economy can affect the return on the investment.

Probability Return

Boom 0.4 25.00%

Good 0.2 15.00%

Level    0.2 10.00%

Slump  0.2 -5.00%

Use the table of returns and probabilities above to determine the expected return on Linda’s investment? (Round answer to 3 decimal places, e.g. 0.076.)

Expected return ______

Use the table of returns and probabilities above to determine the standard deviation of the return on Linda's investment? (Round answer to 5 decimal places, e.g. 0.07680.)

Standard deviation _____
Answer #1

Expected Return on Linda's Investment =0.4*25%+0.2*15%+0.2*10%+0.2*-5% =14.00%or 0.140
Standard Deviation of Linda's Investment =(0.4*(25%-14%)^2+0.2*(15%-14%)^2+0.2*(10%-14%)^2+0.2*(-5%-14%)^2)^0.5 =0.11136